Gerhan Company's flexible budget for the units manufactured in May shows $15,640 of total factory overhead; this output level represents 70% of available capacity. During May, the company applied overhead to production at the rate of $3.00 per direct labor hour (DLH), based on a denominator volume level of 6,120 DLHs, which represents 90% of available capacity. The company used 5,000 DLHs and incurred $16,500 of total factory overhead cost during May, including $6,800 for fixed factory overhead.

What is the factory overhead efficiency variance (to the nearest whole dollar) for May under the assumption that Gerhan uses a four-variance breakdown (decomposition) of the total overhead variance?

Multiple Choice

  • $180 unfavorable.

  • $380 favorable.

  • $380 unfavorable.

  • $480 unfavorable.

  • $480 favorable.
Answer #1

Solution:

Budgeted overhead at 90% capacity = 6120*3 = $18,360

Budgeted overhead at 70% capacity = $15,640

DLH at 70% capacity = 6120/90%*70% = 4760 hours

Variable overhead rate = ($18,360 - $15,640) / (6120 - 4760) = $2 per DLH

actual hours in may = 5000 DLH

Variable overhead efficiency variance = (SH - AH) * SR = (4760 - 5000) * $2 = $480 unfavorable

Hence 4th option is correct.